OpenStudy (emandem):

Use the arithmetic explicit formula to find 11th term of the sequence When F(1)=16 and d=2

OpenStudy (welshfella):

the nth term an = F1 + d(n-1) = 16 + 2(n-1)

OpenStudy (welshfella):

Just plug in n=11 into the above formula

OpenStudy (nkemdirim):